Creative Technologies Corp. has the following information:
Annual sales RM1,200,000
Current liabilities RM375,000
Days sales outstanding (DSO) (365-day year) 40
Inventory Turnover ratio 4.8
Current ratio 1.2
The company’s current assets consist of cash, inventories, and accounts receivable.
How much cash does the company have on its balance sheet?
1
Expert's answer
2016-02-15T00:00:56-0500
By definition, Current Ratio, DSO, ITR are:

Current Ratio = Current Assets / Current Liabilities
therefore
Current Assets = Current Liabilities * Current Ratio = RM375,000 * 1.2 = RM450,000

Days Sales Outstanding = (Accounts Receivable / Annual Sales) * 365
therefore
Accounts Receivable = DSO * Annual Sales / 365 = 40 * RM1,200,000 / 365 = RM131,507

Inventory Turnover Ratio = Annual Sales / Average Inventory
therefore
Average Inventory = Annual Sales / Inventory Turnover Ratio = RM1,200,000 / 4.8 = RM250,000

According to the assignment:
Current Assets = Cash + Inventories + Accounts Receivable
therefore
Cash = Current Assets - Inventories - Account Receivable = RM450,000 - RM250,000 - RM131,507 = RM68,493


The answer is:

The company has RM68,493 cash on its balance sheet